A visitor to a lighthouse wishes to determine the height of the tower. She ties a spool of thread to a small rock to make a simple pendulum, which she hangs down the center of a spiral staircase of the tower. The period of oscillation is 9.67 s. What is the height of the tower?
Answer #1

Solution) T = 9.67 s

h = ?

For simple pendulum

T = (2(pi))(h/g)^(1/2)

Squaring on both sides

T^2 = (4(pi^2))(h/g)

h = ((g)(T^2))/(4(pi^2))

h = ((9.8)(9.67^2))/(4(pi^2))

h = 23.2 m